# Constraining the Primoridal Black Hole Abundance with Space-Based Detectors

### v1

@@ 5 Constraint on PBH Abundance
In this section, we discuss the constraints that various space-based detectors can place on the PBH abundance across different mass ranges. As shown in the previous section, these detectors can constrain the SGWB induced by the curvature perturbation in different frequencies, thereby limiting the amplitude of the curvature perturbation at different wavenumbers. Consequently, these constraints translate into bounds on the existence of PBHs with masses approximately corresponding to the horizon mass at the time when these modes re-enter the horizon.
Figure 4: The constraints of detectable region on amplitude $\mathcal{A}_{\zeta}$ and peak wavenumber $k_{p}$ for power spectra with $\Delta=0$ (upper left), $\Delta=0.28$ (upper right), $\Delta=0.5$ (lower left), $\Delta=1$ (lower right). The solid curves represent the sensitivity of LISA (red), Taiji (green) and TianQin (purple), with a total duration 3 yr and SNR $\varrho=3$. The gray dashed line represents $f_{\rm PBH}=1$.
The range of detectable wavenumbers of the curvature perturbations depends not only on the peak amplitude of the SIGW $\Omega_{\rm GW}(k_{p})\propto\mathcal{A}_{\zeta}^{2}$, but also on its ultraviolet (UV) and infrared (IR) tails. If the curvature perturbation spectrum is monochromatic, the IR tail of SIGW spectrum is $k^{2}$ [88]. As the width of the curvature perturbation spectrum $\Delta$ increases, the infrared spectrum has a broken power-law, from $k^{3}$ ($k\lesssim k_{p}\Delta$) in the far-IR to $k^{2}$ ($k\gtrsim k_{p}\Delta$) in the near-IR [84], which approaches a universal $k^{3}$ scaling as $\Delta\to\mathcal{O}(1)$ [110]. Therefore, a spectrum peaked at frequencies higher than millihertz band is still possible to be probed, thanks to such IR tails. The detectable $A_{\zeta}(k)$ is naturally stretched to the higher wavenumbers, which extends the detectable PBH mass range, covering the aforementioned “mass gap” of $10^{-16}$$10^{-13}~M_{\odot}$. On the other hand, according to $\Omega_{\mathrm{GW}}\propto\mathcal{P}_{\zeta}^{2}$ and (2.11), the UV tail of the SIGW spectrum is also an exponential decay, with a smaller width $\Delta/\sqrt{2}$ [84]. As a result, the lower bound of the detectable wavenumber in $A_{\zeta}(k)$ is nearly a sharp cutoff, which makes the upper bound of the detectable PBH mass $\sim 10^{-11}~M_{\odot}$ robust, which is consistent with the rough estimation from the simple frequency-mass relation (1.1). These features can be clearly seen in Fig. 4 and Fig. 5.
Next we convert the constraints on $A_{\zeta}(k)$ to those on the PBH abundance $f_{\mathrm{PBM}}(M_{\mathrm{PBH}})$. Using the formalism reviewed in Sec. 2, the wavenumber can be mapped to the mass of PBH, while the SIGW amplitude determines the PBH abundance. The extended PBH mass function, illustrated in Fig. 1, can be normalized as
@@
$f_{\rm PBH}$ $=\int{\rm d\,ln}M_{\rm PBH}~f_{\rm PBH}(M_{\rm PBH}),$ (5.2)
$\langle M_{\rm PBH}\rangle$ $=f_{\rm PBH}\left(\int\frac{{\rm d\,ln}M_{\rm PBH}}{M_{\rm PBH}}f_{\rm PBH}(M_{\rm PBH})\right)^{-1}.$ (5.3)
@@
Figure 5: Constraints from different observations on the faction of PBH in DM, $f_{\rm PBH}$, as a function of the PBH average mass $\langle M_{\rm PBH}\rangle$, for PBH mass function showed in Fig. 1 with monochromatic spectrum $\Delta=0$ (upper left) and log-normal spectrum $\Delta=0.28$ (upper right), $\Delta=0.5$ (lower left) and $\Delta=1$ (lower right). The orange region on the left is excluded by evaporations of CMB anisotropies [112] and extragalactic $\gamma$-ray [113]. The right blue region by microlensing (blue shades) of Subaru HSC [9], EROS [114], OGLE [10]. The region between colored dashed lines is the detectable parameters space of LISA (red), Taiji (green) and TianQin (purple), with a total duration 3 yr and SNR $\varrho=3$.
Figure 6: The contour plots of the SNR for induced GW detected by the LISA (left), Taiji (middle), and TianQin (right), corresponding to different widths of the curvature perturbation spectra $\Delta$ (0, 0.28, 0.5 and 1). The black solid lines in the upper left and right corners correspond to the same constraints in Fig.5. The contours are shown in shades of blue, with increasing color intensity indicating larger signal-to-noise ratios. From outer to inner, the contours correspond to $\mathrm{SNR}=3,\,10,\,10^{2},\,10^{3},$ and $10^{4}$, respectively.
@@
$\int{\rm d\,ln}M_{\rm PBH}\,\frac{f_{\rm PBH}F(M_{\rm PBH}/\langle M_{\rm PBH}\rangle)}{f_{\rm mo}(M_{\rm PBH})}\leqslant 1\,$, (5.4)
which ensures that the total contribution from PBHs across all masses does not exceed the observational limit. We note that the constraints from microlensing and evaporation remain nearly unchanged, since they depend only on the shape of the normalized mass function $F_{\Delta}$, which varies only mildly, as shown in Fig. 1. As we mentioned before, for these power spectra, all three interferometers LISA, Taiji, and TianQin can cover the entire PBH-dark-matter mass window sandwiched between microlensing and evaporation constraints. Furthermore, we notice that the lower bounds from space-based detectors, given by IR tails from higher-frequency peaks, depend sensitively on the width of the power spectrum.
@@
$\sum_{j=1}^{N}\left(\int\mathrm{d}\ln M\,\frac{f_{\rm PBH}F(M_{\rm PBH}/\langle M_{\rm PBH}\rangle)}{f_{{\rm mo},j}(M_{\rm PBH})}\right)^{2}\leqslant 1,$ (5.5)
where $f_{{\rm mo},j}(M_{\rm PBH})$ denotes the $j$-th observational constraint. The resulting joint constraints are shown as a grayscale background, with darker regions indicating stronger exclusion. For clarity, we show only the $f_{\rm PBH}=1$ contour from SIGW detection. We notice that the width of the detectable PBH mass range is not a monotonic function of $\Delta$, and there is a minimum at $\Delta\approx 0.28$. This is because, when $\Delta\ll 1$, the IR scaling is $k^{2}$, which allows the ertral peaks to go to higher frequencies. On the other hand when $\Delta\approx\mathcal{O}(1)$, the width of the $\mathcal{P}_{\zeta}$ itself is roughly the width of $\Omega_{\mathrm{GW}}$, which makes the detectable range bigger. We have to emphasize that even for the minimal detectable window with $\Delta\approx 0.28$, it is still sufficient to cover the entire asteroid-mass window, guaranteeing that SIGW associated with PBHs will be detectable if dark matter is composed of asteroid-mass PBHs. Actually, as there is still parameter space beyond the the asteroid-mass window, the SNR of the SIGW signal when PBHs are all the dark matter can be as high as $10^{3}\sim 10^{4}$, as is clearly shown in Fig. 6.
@@
Figure 7: Combined constraints in the $\langle M_{\rm PBH}\rangle$$\Delta$ plane for a log-normal curvature perturbation spectrum with width $\Delta$. The shaded background indicates the current upper limits on the PBH dark-matter fraction $f_{\rm PBH}$ as a function of PBH mass and spectral width $\Delta$, derived from reported evaporation and microlensing constraints for monochromatic spectrum. Darker regions correspond to stronger bounds, with ${\log}_{10}f_{\rm PBH}<-3$ in the black region, while the white region remains weakly constrained by current observations ${\rm log}_{10}f_{\rm PBH}<0$. The colored dashed curves show the sensitivity reaches of space-based gravitational-wave detectors, assuming ${\rm SNR}=3$ and an observation time of $T_{\rm obs}=3{\rm yr}$. The regions enclosed by each pair of dashed curves represent the PBH mass ranges that can be probed under the assumption $f_{\rm PBH}=1$ for a given $\Delta$. Green, red, and purple curves correspond to Taiji, LISA, and TianQin, respectively.
6 Conclusion & Discusion
@@
Probing PBHs as a dark matter candidate is among the key scientific objectives of future space-based gravitational-wave interferometers such as LISA, Taiji, and TianQin. In this work, we have quantitatively assessed the capability of these detectors to constrain the PBH abundance over a broad mass range centered on the asteroid-mass window, which remains largely unconstrained by existing observations. Taking into account the dependence on the shape of the primordial curvature power spectrum, we find that all the space-based interferometers considered here are expected to detect the associated induced gravitatinal-wave signal with signal-to-noise ratios of order $10^{3}\sim 10^{4}$ across the entire mass window, assuming that PBHs constitute all of the dark matter, and largely independent of the specific form of the power spectrum.
@@
The amplitude of SIGWs to be probed is normalized by the abundance of PBHs, which depend on an accurate calculation of their abundance. Two main approaches are widely used in the literature: the peak theory and the Press–Schechter formalism, with the former systematically predicts a higher abundance [61, 59]. In this work, we adopt the peak profile from peak theory and compute the abundance using the Press–Schechter method, following the prescription of [116, 61, 85, 86], as we prefer to avoid subtle issues like properly treating sub-Jeans scale suppression due to radiation pressure, its impact on the resulting PBH mass, the consistent summation of contributions from PBHs formed at different times, etc. Such uncertainties can affect both the overall PBH abundance and the dependence of the mass function on the shape of the primordial power spectrum. Nevertheless, we expect the qualitative features of our results to remain robust. Notably, numerical simulations suggest that the parameter $K$ in Eq. (2.9), which relates the PBH mass to the horizon mass, depends on the pressure gradient determined by the peak profile, and often uses the horizon radius $r_{m}$ instead of the Jeans radius $r_{\rm w}$ [81, 117]. Compared to the fiducial choice $K=1$ adopted in this work, a more accurate treatment may yield PBH masses up to an order of magnitude larger, potentially shifting the low-mass edge of detectable region beyond constraints of Hawking radiation. However, this gap is expected to be closed soon as more X-ray and gamma-ray data are considered from PBH evaporation [118, 119, 120, 121, 122]. For simplicity, we have also neglected the effects of accretion and mergers after PBH formation, which may further modify the mass distribution. More accurate quantitative predictions will require further theoretical and numerical developments.
@@
It is worth emphasizing that the PBH abundance is highly sensitive to the statistical properties of overdense regions, which may be affected by both the non-Gaussianity of curvature perturbations and the statistical formalism used to evaluate the number density and the profiles of peaks. Non-Gaussian features naturally arise in many inflation models that can enhance the power spectrum, such as ultra-slow-roll inflation [123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 130, 133, 134, 72], constant-roll inflation [135, 136, 136, 136, 137, 138, 139, 140, 70], the curvaton scenario [141, 142, 143, 49, 144], stochastic inflaion [145, 146, 147, 148, 149, 150, 151, 152, 153], etc. For instance, if non-Gaussian curvature perturbation can be expanded perturbatively, at quadratic order, we have $\zeta=\zeta_{\rm G}+(3/5)f_{\mathrm{NL}}\zeta_{\rm G}^{2}$. A positive $f_{\mathrm{NL}}$ enhances the PBH formation, which requires smaller curvature perturbation to generate the same amount of PBHs, thus predicts smaller SIGW signal. Up to quadratic level, such a suppression is mild, and PBH dark matter centered in the asteroid-mass window can still be probed by LISA, Taiji, and TianQin with a smaller signal-to-noise ratio [51, 154]. However, the low-mass edge of the window might be undetectable when $f_{\mathrm{NL}}$ is large [155]. For logarithmic non-Gaussianity of the form $\zeta=-(1/\gamma)\ln(1-\gamma\zeta_{\mathrm{G}})$ with $\gamma\lesssim 3.1$ (corresponding to a nominal $f_{\mathrm{NL}}\lesssim 2.6$), the suppression of the SIGW signal is still mild [156, 140]. For larger values of $\gamma$, however, PBHs formed through the bubble channel become dominant [157, 72], thereby leading to a further suppression of the associated SIGW signal. A detailed investigation of this regime is therefore left for future work.
@@
Note Added: While this paper was in preparation, [155] appeared on arXiv, which studies how to probe PBH dark matter with LISA. The local non-Gaussianity of the curvature perturbation is also included. Our result for LISA in the Gaussian case is in agreement with [155], and our paper also includes the analysis of Taiji and TianQin.
Acknowledgement

### v2

@@ 5 Constraint on PBH Abundance
In this section, we discuss the constraints that various space-based detectors can place on the PBH abundance across different mass ranges. As shown in the previous section, these detectors can constrain the SGWB induced by the curvature perturbation at different frequencies, thereby limiting the amplitude of the curvature perturbation at different wavenumbers.  [^2] Primordial black hole binaries also emit GWs [152, 35, 146, 147], which form a SGWB by their superposition [8, 7, 115, 168, 32]. The peak of such a SGWB from asteroid-mass PBH binaries is higher than GHz, which is difficult to be probed and is not discussed in this paper. Consequently, these constraints translate into bounds on the existence of PBHs with masses approximately corresponding to the horizon mass at the time when these modes re-enter the horizon.
Figure 4: The constraints on the detectable region in amplitude $\mathcal{A}_{\zeta}$ and peak wavenumber $k_{p}$ for power spectra with $\Delta=0$ (upper left), $\Delta=0.28$ (upper right), $\Delta=0.5$ (lower left), $\Delta=1$ (lower right). The solid curves represent the sensitivity of LISA (red), Taiji (green) and TianQin (purple), with a total duration of $3\,\mathrm{yr}$ and SNR $\varrho=3$. The gray dashed line represents $f_{\rm PBH}=1$.
The range of detectable wavenumbers of the curvature perturbations depends not only on the peak amplitude of the SIGW $\Omega_{\rm GW}(k_{p})\propto\mathcal{A}_{\zeta}^{2}$, but also on its ultraviolet (UV) and infrared (IR) tails. If the curvature perturbation spectrum is monochromatic, the IR tail of the SIGW spectrum is $k^{2}$ [103]. As the width of the curvature perturbation spectrum $\Delta$ increases, the infrared spectrum has a broken power-law, from $k^{3}$ ($k\lesssim k_{p}\Delta$) in the far-IR to $k^{2}$ ($k\gtrsim k_{p}\Delta$) in the near-IR [139], which approaches a universal $k^{3}$ scaling as $\Delta\to\mathcal{O}(1)$ [39]. Therefore, a spectrum peaked at frequencies higher than millihertz band is still possible to be probed, thanks to such IR tails. The detectable $A_{\zeta}(k)$ is naturally stretched to the higher wavenumbers, which extends the detectable PBH mass range, covering the aforementioned “mass gap” of $10^{-16}$$10^{-13}~M_{\odot}$. On the other hand, according to $\Omega_{\mathrm{GW}}\propto\mathcal{P}_{\zeta}^{2}$ and (2.11), the UV tail of the SIGW spectrum is also an exponential decay, with a smaller width $\Delta/\sqrt{2}$ [139]. As a result, the lower bound of the detectable wavenumber in $A_{\zeta}(k)$ is nearly a sharp cutoff, and thereby makes the upper bound of the detectable PBH mass $\sim 10^{-11}\,M_{\odot}$ robust, which is consistent with the rough estimation from the simple frequency-mass relation (1.1). These features can be clearly seen in Fig. 4 and Fig. 5.
Next we convert the constraints on $A_{\zeta}(k)$ to those on the PBH abundance $f_{\mathrm{PBH}}(M_{\mathrm{PBH}})$. Using the formalism reviewed in Sec. 2, the wavenumber can be mapped to the mass of PBH, while the SIGW amplitude determines the PBH abundance. The extended PBH mass function, illustrated in Fig. 1, can be normalized as
@@
$f_{\rm PBH}$ $=\int{\rm d}\ln M_{\rm PBH}~f_{\rm PBH}(M_{\rm PBH}),$ (5.2)
$\langle M_{\rm PBH}\rangle$ $=f_{\rm PBH}\left(\int\frac{{\rm d}\ln M_{\rm PBH}}{M_{\rm PBH}}f_{\rm PBH}(M_{\rm PBH})\right)^{-1}.$ (5.3)
@@
Figure 5: Constraints from different observations on the fraction of PBH in DM, $f_{\rm PBH}$, as a function of the PBH average mass $\langle M_{\rm PBH}\rangle$, for PBH mass function shown in Fig. 1 with monochromatic spectrum $\Delta=0$ (upper left) and log-normal spectrum $\Delta=0.28$ (upper right), $\Delta=0.5$ (lower left) and $\Delta=1$ (lower right). The orange region on the left is excluded by constraints from evaporation effects on CMB anisotropies [4] and extragalactic $\gamma$-rays [48]. The right blue region is excluded by microlensing (blue shades) of Subaru HSC [135], EROS [161], and OGLE [124]. The region between colored dashed lines is the detectable parameter space of LISA (red), Taiji (green), and TianQin (purple), with a total duration of $3\mathrm{yr}$ and SNR $\varrho=3$.
Figure 6: The contour plots of the SNR for induced GWs detected by the LISA (left), Taiji (middle), and TianQin (right), corresponding to different widths of the curvature perturbation spectra $\Delta$ (0, 0.28, 0.5 and 1). The black solid lines in the upper left and right corners correspond to the same constraints in Fig. 5. The contours are shown in shades of blue, with increasing color intensity indicating larger signal-to-noise ratios. From outer to inner, the contours correspond to $\mathrm{SNR}=3,\,10,\,10^{2},\,10^{3},$ and $10^{4}$, respectively.
@@
$\int{\rm d}\ln M_{\rm PBH}\,\frac{f_{\rm PBH}F(M_{\rm PBH}/\langle M_{\rm PBH}\rangle)}{f_{\rm mo}(M_{\rm PBH})}\leqslant 1\,$, (5.4)
which ensures that the total contribution from PBHs across all masses does not exceed the observational limit. We note that the constraints from microlensing and evaporation remain nearly unchanged, since they depend only on the shape of the normalized mass function $F_{\Delta}$, which varies only mildly, as shown in Fig. 1. As mentioned before, for these power spectra, all three interferometers LISA, Taiji, and TianQin can cover the entire PBH-dark-matter mass window sandwiched between microlensing and evaporation constraints. Furthermore, we notice that the lower bounds from space-based detectors, given by IR tails from higher-frequency peaks, depend sensitively on the width of the power spectrum.
@@
$\sum_{j=1}^{N}\left(\int\mathrm{d}\ln M_{\rm PBH}\,\frac{f_{\rm PBH}F(M_{\rm PBH}/\langle M_{\rm PBH}\rangle)}{f_{{\rm mo},j}(M_{\rm PBH})}\right)^{2}\leqslant 1,$ (5.5)
where $f_{{\rm mo},j}(M_{\rm PBH})$ denotes the $j$-th observational constraint. The resulting joint constraints are shown as a grayscale background, with darker regions indicating stronger exclusion. For clarity, we show only the $f_{\rm PBH}=1$ contour from SIGW detection. We notice that the width of the detectable PBH mass range is not a monotonic function of $\Delta$, and there is a minimum at $\Delta\approx 0.28$. This is because, when $\Delta\ll 1$, the IR scaling is $k^{2}$, which allows the GW peak to go to higher frequencies. On the other hand, $\Delta\approx\mathcal{O}(1)$, the width of the $\mathcal{P}_{\zeta}$ itself is roughly the width of $\Omega_{\mathrm{GW}}$, which makes the detectable range bigger. We have to emphasize that even for the minimal detectable window with $\Delta\approx 0.28$, it is still sufficient to cover the entire asteroid-mass window, guaranteeing that SIGW associated with PBHs will be detectable if dark matter is composed of asteroid-mass PBHs. Actually, as there is still parameter space beyond the asteroid-mass window, the SNR of the SIGW signal when PBHs are all the dark matter can be as high as $10^{3}$$10^{4}$, as is clearly shown in Fig. 6.
@@
Figure 7: Combined constraints in the $\langle M_{\rm PBH}\rangle$$\Delta$ plane for a log-normal curvature perturbation spectrum with width $\Delta$. The shaded background indicates the current upper limits on the PBH dark-matter fraction $f_{\rm PBH}$ as a function of PBH mass and spectral width $\Delta$, derived from reported evaporation and microlensing constraints for a monochromatic spectrum. Darker regions correspond to stronger bounds, with ${\log}_{10}f_{\rm PBH}<-3$ in the black region, while the white region remains weakly constrained by current observations ${\log}_{10}f_{\rm PBH}<0$. The colored dashed curves show the sensitivity reaches of space-based gravitational-wave detectors, assuming ${\rm SNR}=3$ and an observation time of $T_{\rm obs}=3\,\mathrm{yr}$. The regions enclosed by each pair of dashed curves represent the PBH mass ranges that can be probed under the assumption $f_{\rm PBH}=1$ for a given $\Delta$. Green, red, and purple curves correspond to Taiji, LISA, and TianQin, respectively.
6 Conclusion and Discussions
@@
Probing PBHs as a dark matter candidate is among the key scientific objectives of future space-based gravitational-wave interferometers such as LISA, Taiji, and TianQin. In this work, we have quantitatively assessed the capability of these detectors to constrain the PBH abundance over a broad mass range centered on the asteroid-mass window, which remains largely unconstrained by existing observations. Taking into account the dependence on the shape of the primordial curvature power spectrum, we find that all space-based interferometers considered here are expected to detect the associated induced gravitational-wave signal with signal-to-noise ratios of order $10^{3}$$10^{4}$ across the entire mass window, assuming that PBHs constitute all of the dark matter, and largely independent of the specific form of the power spectrum.
@@
The amplitude of SIGWs to be probed is normalized by the abundance of PBHs, which depend on an accurate calculation of their abundance. Two main approaches are widely used in the literature: the peak theory and the Press–Schechter formalism, of which the former systematically predicts a higher abundance [102, 138]. In this work, we adopt the peak profile from peak theory and compute the abundance using the Press–Schechter method, following the prescription of [33, 102, 82, 142], as we prefer to avoid subtle issues like properly treating sub-Jeans scale suppression due to radiation pressure, its impact on the resulting PBH mass, the consistent summation of contributions from PBHs formed at different times, etc. Such uncertainties can affect both the overall PBH abundance and the dependence of the mass function on the shape of the primordial power spectrum. Nevertheless, we expect the qualitative features of our results to remain robust. Notably, numerical simulations suggest that the parameter $K$ in Eq. (2.9), which relates the PBH mass to the horizon mass, depends on the pressure gradient determined by the peak profile, and often uses the horizon radius $r_{m}$ instead of the Jeans radius $r_{\rm w}$ [71, 127]. Compared to the fiducial choice $K=1$ adopted in this work, a more accurate treatment may yield PBH masses up to an order of magnitude larger, potentially shifting the low-mass edge of the detectable region beyond the constraints of Hawking radiation. However, this gap is expected to be closed soon as more X-ray and gamma-ray data are considered from PBH evaporation [62, 106, 105, 88, 121]. For simplicity, we have also neglected the effects of accretion and mergers after PBH formation, which may further modify the mass distribution. More accurate quantitative predictions will require further theoretical and numerical developments.
@@
It is worth emphasizing that the PBH abundance is highly sensitive to the statistical properties of overdense regions, which may be affected by both the non-Gaussianity of curvature perturbations and the statistical formalism used to evaluate the number density and the profiles of peaks. Non-Gaussian features naturally arise in many inflation models that can enhance the power spectrum, such as ultra-slow-roll inflation [130, 50, 41, 34, 136, 140, 19, 25, 57, 169, 58, 45, 69], constant-roll inflation [22, 21, 21, 21, 145, 66, 170, 90, 155], the curvaton scenario [153, 141, 77, 79, 87], stochastic inflation [137, 20, 163, 144, 26, 94, 125, 36, 131], etc. For instance, if non-Gaussian curvature perturbation can be expanded perturbatively, at quadratic order, we have $\zeta=\zeta_{\rm G}+(3/5)f_{\mathrm{NL}}\zeta_{\rm G}^{2}$. A positive $f_{\mathrm{NL}}$ enhances the PBH formation, which requires smaller curvature perturbation to generate the same amount of PBHs, thus predicts smaller SIGW signal. Up to quadratic level, such a suppression is mild, and PBH dark matter centered in the asteroid-mass window can still be probed by LISA, Taiji, and TianQin with a smaller signal-to-noise ratio [38, 143]. However, the low-mass edge of the window might be undetectable when $f_{\mathrm{NL}}$ is large [93]. For logarithmic non-Gaussianity of the form $\zeta=-(1/\gamma)\ln(1-\gamma\zeta_{\mathrm{G}})$ with $\gamma\lesssim 3.1$ (corresponding to a nominal $f_{\mathrm{NL}}\lesssim 2.6$), the suppression of the SIGW signal is still mild [3, 90]. For larger values of $\gamma$, however, PBHs formed through the bubble channel become dominant [68, 69], thereby leading to a further suppression of the associated SIGW signal. A detailed investigation of this regime is therefore left for future work.
@@
Note Added: While this paper was in preparation, [93] appeared on arXiv, which studies how to probe PBH dark matter with LISA. The local non-Gaussianity of the curvature perturbation is also taken into account. Our result for LISA in the Gaussian case is in agreement with [93], and our paper also includes the analysis of Taiji and TianQin.
Acknowledgements